# Benchmarking Hindi LLMs: A New Suite of Datasets and a Comparative Analysis

### v1

@@ 1 Introduction
In recent years, significant progress has been made in developing evaluation resources for Indic languages, typified by benchmarks such as IndicGLUE Kakwani et al. (2020), MILU Verma et al. (2025), IndicMMLU-Pro Sankalp et al. , and IndicGenBench Singh et al. (2024a). These resources have been instrumental in assessing the core capabilities of pre-trained base models across numerous languages of the Indian subcontinent. Despite this progress, the existing benchmarks primarily target pre-trained base models, leaving a noticeable gap in resources for assessing the capabilities of instruction-tuned models. Consequently, benchmarks for critical skills like instruction following, conversational ability, and function calling, such as Hindi versions of IFEval, MT-Bench, and BFCL, are largely unavailable publicly.
@@
  • We introduce a suite of five new, high-quality benchmarks (IFEval-Hi, MT-Bench-Hi, GSM8K-Hi, ChatRAG-Hi, and BFCL-Hi) for evaluating instruction-tuned LLMs in Hindi and detail the curation process developed for their creation.
  • We present a comprehensive benchmark of prominent, publicly available LLMs on this new suite, providing the first robust comparative analysis of their capabilities in Hindi. Our findings show that while specialized models exhibit strength in specific tasks, Gemma-2-9b-it in the SLM class and GPT-OSS-120B in the LLM class emerge as the most capable general-purpose models.
@@ 3.3 GSM8K-Hi
Model Size MT-Bench-Hi BFCL-Hi GSM8K-Hi IFEval-Hi ChatRAG-Hi SLMs Gemma-2-2b-it 2B 4.37 32.96 26.99 38.92 29.89 Llama-3.2-3B-Instruct 3B 5.14 33.81 40.11 40.80 32.60 Nemotron-Mini-4B-Instruct 4B 3.44 - 32.22 36.08 27.32 Nemotron-4-Mini-Hindi-4B-Instruct 4B 6.01 52.82 47.31 51.65 36.07 Llama-3.1-8B-Instruct 8B 6.44 31.23 61.33 48.82 38.03 Aya-expanse-8b 8B 6.58 36.56 64.52 42.92 30.15 Gemma-2-9b-it 9B 7.37 50.51 64.44 61.79 40.97 Krutrim-2-instruct 12B 6.31 26.88 56.56 59.32 37.48 LLMs GPT-OSS-20B (reasoning low) 21B 8.51 54.60 80.64 69.04 26.16 Mistral-Small-3.2-24B-Instruct-2506 24B 7.83 41.45 77.55 66.89 37.92 Sarvam-M (reasoning off) 24B 8.25 48.60 82.30 71.64 40.14 Gemma-3-27b-it 27B 8.31 62.42 78.12 67.72 45.23 GPT-OSS-120B (reasoning low) 117B 8.70 61.26 93.41 73.86 29.85 Qwen3-235B-A22B-FP8 (reasoning off) 235B 8.10 59.88 89.69 68.11 32.47 Llama-3.1-405B 405B 7.17 49.53 86.27 68.66 47.46

### v2

@@ 1 Introduction
In recent years, significant progress has been made in developing evaluation resources for Indic languages, typified by benchmarks such as IndicGLUE Kakwani et al. (2020), MILU Verma et al. (2025), IndicMMLU-Pro Sankalp et al. , and IndicGenBench Singh et al. (2024a). These resources have been instrumental in assessing the core capabilities of pre-trained base models across numerous languages of the Indian subcontinent Joshi et al. (2024). Despite this progress, the existing benchmarks primarily target pre-trained base models, leaving a noticeable gap in resources for assessing the capabilities of instruction-tuned models. Consequently, benchmarks for critical skills like instruction following, conversational ability, and function calling, such as Hindi versions of IFEval, MT-Bench, and BFCL, are largely unavailable publicly.
@@ 3.3 GSM8K-Hi
Model Size MT-Bench-Hi BFCL-Hi GSM8K-Hi IFEval-Hi ChatRAG-Hi SLMs Gemma-2-2b-it 2B 4.37 32.96 26.99 38.92 29.89 Llama-3.2-3B-Instruct 3B 5.14 33.81 40.11 40.80 32.60 Nemotron-Mini-4B-Instruct 4B 3.44 - 32.22 36.08 27.32 Nemotron-4-Mini-Hindi-4B-Instruct 4B 6.01 52.82 47.31 51.65 36.07 Llama-3.1-8B-Instruct 8B 6.44 31.23 61.33 48.82 38.03 Aya-expanse-8b 8B 6.58 36.56 64.52 42.92 30.15 Gemma-2-9b-it 9B 7.37 50.51 64.44 61.79 40.97 Krutrim-2-instruct 12B 6.31 26.88 56.56 59.32 37.48 LLMs GPT-OSS-20B (reasoning low) 21B 8.51 54.60 80.64 69.04 26.16 Mistral-Small-3.2-24B-Instruct-2506 24B 7.83 41.45 77.55 66.89 37.92 Sarvam-M (reasoning off) 24B 8.25 48.60 82.30 71.64 40.14 Gemma-3-27b-it 27B 8.31 62.42 78.12 67.72 45.23 GPT-OSS-120B (reasoning low) 117B 8.70 61.26 93.41 73.86 29.85 Qwen3-235B-A22B-FP8 (reasoning off) 235B 8.10 59.88 89.69 68.11 32.47 Llama-3.1-405B 405B 7.17 49.53 86.27 68.66 47.46 LLMs (Reasoning) GPT-OSS-20B (reasoning medium) 21B 8.43 63.26 83.41 72.01 29.16 GPT-OSS-20B (reasoning high) 21B 8.23 64.77 83.44 72.11 32.39 Sarvam-M (reasoning on) 24B 8.60 59.53 84.40 74.06 37.13 GPT-OSS-120B (reasoning medium) 117B 8.79 66.19 95.93 76.69 30.80 GPT-OSS-120B (reasoning high) 117B 8.70 64.90 96.27 76.80 31.82
@@ 4 Results and Discussion
Furthermore, activating the dedicated reasoning modes in models like GPT-OSS and Sarvam-M provides a substantial performance uplift on complex tasks like BFCL, GSM8K, and IFEval. With these capabilities enabled, GPT-OSS-120B achieves the top scores across multiple benchmarks, highlighting the value of reasoning models for Hindi.
@@ 5 Conclusion
Acknowledgements
This work would not have been possible without contributions from many people at NVIDIA. To mention a few: Priyanka Chavan, Shweta Dash, Asmita Kadwe, Nilesh Chauhan, Saloni Pipada, Shrutika Marke, Priyanka Patil, Ninad Nemade, Muskan Grewal, Aditya Patil, and Noopur Mishra.